OpenStudy (anonymous):

what does linear and nonlinear mean? is this problem linear y=0

OpenStudy (anonymous):

A linear equation is of the form \(y=ax+b\). A nonlinear equation contains more complicated functions of \(x\), like \(x^2\) and \(\sin x\).

OpenStudy (anonymous):

Basically linear is straight line while non-linear is curvy :)

OpenStudy (acxbox22):

nonlinear means a function that is not constant y=mxz=b is linear y=x^2, y=4^x, y=sqrt x is nonlinear

OpenStudy (acxbox22):

the second line is supposed to say y=mx+b is linear

OpenStudy (anonymous):

isit yes or no

OpenStudy (anonymous):

YES